OpenStudy (anonymous):

There are (32)4 ⋅ 30 bacteria in a petri dish. What is the total number of bacteria in the dish?

OpenStudy (anonymous):

38 39 311 324

pooja195 (pooja195):

\[\huge~\rm~32 \times 4 =?\]

OpenStudy (anonymous):

no its 3 to power of 2

pooja195 (pooja195):

Oh ok \[\huge~\rm~3^2= ? \]

OpenStudy (anonymous):

9 ?

pooja195 (pooja195):

Something isn't right....

pooja195 (pooja195):

(3^2)(4)(30) =(9)(4)(30) =(36)(30) =1080 This isn't one of your choices ;-;

OpenStudy (anonymous):

no

pooja195 (pooja195):

@TheSmartOne ? Can you help?

OpenStudy (anonymous):

it 3 to the power of 2 in () then power of 4 * 3 to the power 0

pooja195 (pooja195):

Can you screenshot the problem?

OpenStudy (anonymous):

there it is

pooja195 (pooja195):

(32)4(30) =(94)(30) =6561(30) =(6561)(1) =6561

pooja195 (pooja195):

Which of your options is equivalent to that?

OpenStudy (anonymous):

A?